# A Quality-Centric Framework for Generic Deepfake Detection

### v2

@@ Abstract
This paper addresses the generalization issue in deepfake detection by harnessing forgery quality in training data. Generally, the forgery quality of different deepfakes varies: some have easily recognizable forgery clues, while others are highly realistic. Existing works often train detectors on a mix of deepfakes with varying forgery qualities, potentially leading detectors to short-cut the easy-to-spot artifacts from low-quality forgery samples, thereby hurting generalization performance. To tackle this issue, we propose a novel quality-centric framework for generic deepfake detection, which is composed of a Quality Evaluator, a low-quality data enhancement module, and a learning pacing strategy that explicitly incorporates forgery quality into the training process. The framework is inspired by curriculum learning, which is designed to gradually enable the detector to learn more challenging deepfake samples, starting with easier samples and progressing to more realistic ones. We employ both static and dynamic assessments to assess the forgery quality, combining their scores to produce a final rating for each training sample. The rating score guides the selection of deepfake samples for training, with higher-rated samples having a higher probability of being chosen. Furthermore, we propose a novel frequency data augmentation method specifically designed for low-quality forgery samples, which helps to reduce obvious forgery traces and improve their overall realism. Extensive experiments show that our method can be applied in a plug-and-play manner and significantly enhance the generalization performance.
Index Terms:
Deepfake Detection, Multimedia Forensics, Forgery Quality Evaluation
I Introduction
@@
Deepfake techniques have experienced rapid advancements in recent years, providing a new way of entertainment. Unfortunately, this technology is also broadly misused to spread false misinformation and even bring political influence. To prevent the abuse of deepfake, it is increasingly crucial to develop a reliable deepfake detector. Most earlier deepfake detectors [1, 2, 3, 4, 5, 6] achieve satisfied and promising detection results in the within-domain evaluation scenario, where the distributions of the training and testing data are similar. However, when encountering cross-domain evaluation, with previously unseen forgery methods or data sources, the performance of these models will drop significantly [7]. Thus, the generalization ability of these prior works is still limited.
Figure 1: Illustration of the proposed forgery quality assessment. We assess the forgery quality from two views: swapping pairs (static) and model feedback (dynamic). Our quality-centric framework enables the detector to progressively learn the deepfake artifacts from easy (lower quality) to hard (higher quality), avoiding the potential shortcuts on easy artifacts from low-quality samples, thereby improving the generalization.
Figure 2: Comparison of generalization performance of models trained using data of different quality. Baseline refers to using all data to train the model, LQ only means that it only uses low static quality data for training, and HQ stands for training using high static quality data.
@@
One reason for the generalization issue is the model’s shortcuts [8] on the easy-to-spot artifacts from the training samples with low forgery quality [9, 10, 11]. For instance, the “irregular swap” fake samples are frequently observed in FF++ (a common training set) but rarely in CDF-v2 (a common testing set). “Irregular swaps” is firstly noticed by [11], where face-swapping involves pairs with distinct facial attributes, such as differences in race or gender. These irregular swaps often severely distort the human facial structure, leading to lower-quality fake samples. As illustrated in Figure 1, we see that when “irregular swaps” occur (i.e., swapping a white face with a black face or a female face with a male face), the easy-to-recognized deepfake samples are generated. To verify the impact of these lower forgery quality fakes on the model’s generalization, we selectively choose samples with different forgery qualities for training the same detector (i.e., Xception [12]). Results in Figure 2 verify that these low static quality samples indeed hurt the model’s generalization, especially in the Wild [13] and DFDC-P [14] datasets, where simply training with higher-quality samples can obtain about 10% improvement in the generalization performance. However, most existing methods often train detectors on a mix of qualities of deepfake data, which might inevitably struggle with this issue (similar to the “baseline” in Figure 2), as they treat all data as the same and ignore the impact of different forgery qualities.
@@
These observations motivate us to treat training data with forgery qualities differently. To this end, we need to address three critical considerations first: Question 1: How can we assess the quality level of a sample? Question 2: How to deal with low-quality fake data? Question 3: How can we maximize the utilization of all types of data? In the following content, we will answer the above research questions in detail.
@@
For Question 1, our solution is to assess the forgery quality from the perspective of data (in a statical view) and that of a detection model (in a dynamical view), as illustrated in Figure 1. From the static view, we concentrate on swapping pairs in the face-swapping process, with the assumption that obtaining paired data is feasible. For a given deepfake generator, similar IDs for swapping can potentially yield more realistic deepfakes (see Fig. 1 for illustration). The rationale behind this is that: for a high static quality deepfake sample, the visual forgery artifacts should be as small as possible, meaning that the facial structure of the face should be natural and seamless. On the contrary, the low static quality sample contains obvious visual artifacts, leading to significant damage to the facial structure. Therefore, we can assess the forgery quality statically by focusing on the swapping pairs. To achieve this, we employ the face recognition network (i.e., ArcFace [15]), taking both the fake and the corresponding real samples as input (with identical backgrounds), and then compute the cosine similarity of the obtained features. If the similarity is low, we can argue that the fake should be of low static quality. From the dynamic view, we assess the forgery quality from the model’s feedback on each sample during training; a lower prediction probability indicates a higher dynamic score, as the fakes are more challenging to detect (see Figure  1 for illustration). Finally, we combine both the static and dynamic assessment results to produce the Forgery Quality Score (FQS) for sample selection. Note that, unless otherwise specified, low-quality and high-quality data refer to samples characterized by low FQS and high FQS, respectively.
For Question 2, our solution is to reduce the obvious forgery traces of the low-quality samples and enhance their realism, rather than directly discarding them. To achieve this, we propose Frequency Data Augmentation (FreDA), which retains the low-frequency features from real faces and reconstructs new augmented samples by integrating the high-frequency features from the original fake samples. Consequently, the facial structure of augmented low-quality samples is improved, transforming them into relatively high-quality samples. FreDA operates based on the fact that low-frequency components in an image represent the overall facial structure, which are crucial for maintaining the face’s semantic integrity. Since low-frequency regions dominate a fake face and can vary significantly between samples, detectors trained on these features often learn non-generalizable artifacts. In contrast, high-frequency components capture finer details. They are more consistent and robust forgery indicators. FreDA aims to combine these high-frequency parts (more general) of the fake and the low-frequency parts (semantic facial structure) of the real, thereby resulting in a new (more realistic) fake.
Figure 3: Illustration of fake examples with different static quality within the FF++ dataset (training set). Intuitively, training with low static quality samples can lead the model to learn the shortcut and thus hurt the model’s generalization.
For Question 3, our solution is to employ the curriculum learning technology to guide the model to start with easy samples and then gradually focus more on challenging samples. Specifically, we compute the FQS values for each sample and use this value to sample training data during the training process. The higher the FQS score, the more likely it is to be sampled into the hard sample pool. As training progresses, we gradually reduce the number of training samples to focus the model’s attention on harder samples, guiding the progression of the training subsets from easy to hard. Concurrently, we apply FreDA operations to low-quality samples to increase their hardness and enhance the diversity of the training data. In this manner, we can fully leverage training data with all qualities. Experimental results demonstrate that our method can improve both the within- and cross-dataset performance of various kinds of deepfake detectors in a plug-and-play manner.
Compared with our preliminary work [16], this paper mainly makes two additional major contributions. Firstly, the focus is different. [16] mainly concentrates on image quality, whereas our work is more centered on forgery quality. Specifically, the key technical difference lies in the design of the static prior knowledge part. [16] employs the Image Forgery Quality Assessment network (IFQA) to consider the impact of post-processing on image quality but fails to take into account irregular face-swapping scenarios (forgery quality). In this paper, we incorporate the feature similarity measured by ArcFace to evaluate the quality of swapped faces. By making use of the ArcFace score, we can refine the selection of hard samples more accurately, considering samples with better-swapped faces as more challenging ones. Secondly, we have also enhanced how we handle easy samples. While DFFC adopts common data post-processing methods such as JPEG compression and Gaussian blurring for easy samples, we devise a novel data augmentation approach called FreDA, explicitly leveraging frequency domain processing, effectively enhancing all types of forged samples (FF++, BI-based, and SBI-based); therefore it further boosts the generalization capability. Notably, our method outperforms the baseline [12, 17, 18] by about 10% on average across several widely-used evaluation datasets, such as Celeb-DF [19] and DFDC [20].
Our main contributions are as follows:
@@
  • We propose an assessment method combining both static and dynamic views to obatin the Forgery Quality Score (FQS), which allows us to rank the hardness of the training samples, thereby implementing the training sample selection.
  • For low-quality samples, we propose Frequency Data Augmentation (FreDA) to reduce the obvious forgery traces and thereby enhance the realism of the augmented low-quality data, rather than directly discarding them. FreDA can reduce the traces of forgery in forged samples. Using samples processed by FreDA enhances the model’s generalization.
@@
II Related Works
II-A Deepfake Detection
@@
Detecting deepfake videos has become an area of intense research due to the growing concern over the potential misuse of manipulated media content. Most works at an early stage, mainly focus on hand-crafted features, e.g., eye-blinking [21], inconsistencies [22] of head poses [23], and other visual biological artifacts. With the rapid development of deep learning, data-driven-based detectors [2] have shown better performance than the conventional hand-crafted approaches. Zhao et al. [24] used attention mechanisms to enhance the manipulated trace in deepfake videos. Yet, these approaches often suffer from poor generalization when there is a distribution shift between training and testing forgeries [25, 26, 27].
@@
Existing works have sought to improve generalization from different directions. Below, we give a brief introduction by focusing on three key aspects:
@@
1) Forgery Artifact Learning: A key challenge in deepfake detection is learning universal forgery traces that can be generalized across various manipulation methods. Several solutions, such as disentanglement learning [28, 29], reconstruction learning [30, 31] and inconsistent learning [32, 33, 34, 35, 36, 37], have been proposed to learn general forgery artifacts. These methods aim to isolate the common characteristics of manipulated content, irrespective of the specific forgery technique employed. To illustrate, UCF [28] proposed a disentanglement-based framework to eliminate the overfitting of both forgery-irrelevant content and forgery-specific artifacts. UIA-ViT [32] used the self-attention mechanism to make the attention map between patch embeddings naturally represent the consistency relationship. FoCus [34] located forgery cues in unpaired faces and generated more accurate forgery images through a classification-attention region proposal module and a complementary learning module. CADDM [37] used the identity inconsistency between the inner and outer faces for detection. Additionally, RECCE [30] proposed a reconstruction-based framework that performs self-reconstruction using only genuine samples for the learning of common forgery artifacts.
However, these approaches mostly operate at the spatial level but overlook frequency-level artifacts, which have been demonstrated to be crucial for detection [38]. To this end, several studies have incorporated frequency information to enhance the performance of detectors [39, 38, 40, 41, 42]. For example, Ricard et al.[39] show that convolution-based upsampling methods, commonly used in deepfake technologies, can result in a discrepancy in the spectral distribution between fake and real images and videos. To tackle this issue, the authors introduce a frequency-based approach that outperforms many widely used spatial-based methods. Moreover, learning-based frequency domain methods have also been extensively investigated. For instance, Qianet al.[38] develop a set of learnable filters to adaptively mine frequency forgery clues using frequency-aware image decomposition. Additionally, Liuet al. [40] utilize frequency domain information as the fourth channel, combining it with RGB domain information as input to the network. SFDG [41] proposed a space-frequency dynamic graph method to mine relational perception features in the spatial and frequency domains through dynamic graph learning. CD-NeT [42] proposed a cross-domain face forgery detection framework that exploits the consistency of multi-frame correlation representation and complementary cues in RGB and frequency domains.
Additionally, several recent works [43, 44, 45] have proposed to leverage both local and global interactions to more accurately capture forgery traces. DL [43] proposed to use segmentation maps to provide high-level semantic information clues of the image, combine them with noise maps to capture low-level clues, and finally combine the features of the two to distinguish fake faces. MSVT [44] combined the local and global features to mine more detailed spatial-temporal information and integrate multi-level features through the global-local Transformer. Furthermore, [46, 47, 48] leveraged contrastive learning to improve the detection performance of deepfake detectors.
@@
2) Data-level Augmentation: Forgery augmentation techniques have proven to be highly effective in enhancing the generalization ability of deepfake detectors by encouraging the model to learn generic forgery artifacts. Pioneering works such as BI (Face X-Ray) [7] synthesized blended faces by replicating blending artifacts between pairs of pristine images with similar facial landmarks. I2G [49] employed a similar strategy, leveraging pair-wise self-consistency learning to identify inconsistencies within generated fake images. Later efforts, like SBI [10], focused on self-blending, selecting the same source and target faces to train models, while SLADD [50] utilized adversarial training to generate more challenging blending scenarios. AUNet [51] built relationships between action units and used tampered action unit prediction modules to generate feature-level pseudo samples. These approaches highlight the importance of augmenting data to expose models to diverse forgery examples. More recently, [52] and [53] have explored video-level blending for data augmentation. While these approaches have improved generalization, many methods, particularly those that focus on low-level blending artifacts, remain vulnerable to post-processing operations such as compression. For example, Yan et al.[9] demonstrated that compression could significantly degrade the performance of RGB-based augmentation methods like SBI [10] and Face X-Ray [7], highlighting the need for more robust detection frameworks.
@@
3) Quality-aware deepfake detection: Another notable direction is to train the detector with different image qualities (combining both low quality compression images and high quality non-compression ones), aiming to learn the quality-irrelevant forgery artifacts [54]. Recent works [55, 16] proposed to investigate the correlation between the detection performance and the image quality. Additionally, [56] proposed a supper-resolution model for improving low quality deepfake detection. However, most existing works define the “quality” as the image quality, e.g., whether the image is blurry or not. Although our framework is also quality-based, our focus is more on the forgery quality, i.e., whether the deepfake image is realistic or not, rather than the image quality. In our framework, we propose to assess the forgery quality from two distinct aspects: swapping pairs and model feedback. We then design a curriculum-learning-based method to leverage samples with varying forgery qualities dynamically for training, guiding the detector to learn the deepfake samples from easy to hard gradually.
II-B Curriculum Learning
Curriculum Learning (CL) is a training paradigm that seeks to improve the performance of deep neural networks by presenting training data in a meaningful order or curriculum. In CL, training data is divided into several stages, each stage representing a different level of hardness. This approach allows the model to first learn easier examples and then gradually increasing the hardness of the examples. Early works on curriculum learning such as BabyStep [57, 58] have shown that providing the training proceeding with an optimized order of training sets can improve models’ performance. Subsequently, CL has been extended to incorporate different strategies such as dynamic curriculum learning [59] and self-paced learning [60]. Dynamic curriculum learning adjusts the hardness of examples dynamically during training based on the model’s performance, while self-paced learning encourages the model to learn from easier examples first and then gradually move to more complex examples. Self-Paced Learning (SPL) selects samples with smaller loss, and gradually increases the subset size over time to cover all the training data. Kong et al. [61] used a pre-trained model and the current model to calculate the loss of training samples as a definition of hardness. Furthermore, several studies have investigated the effectiveness of CL in various domains, including computer vision and natural language processing. Our method introduces a FQS inspired by the concept of SPL, selecting samples with higher FQS, and gradually decreasing the subset size over time to make the model focus on hard samples.

### v3

# Not All Fakes are Equal: A Quality-Centric Framework for Deepfake Detection

@@ Abstract
Detecting AI-generated images, particularly deepfakes, has become increasingly crucial, with the primary challenge being the generalization to previously unseen manipulation methods. This paper tackles this issue by leveraging the forgery quality of training data to improve the generalization performance of existing deepfake detectors. Generally, the forgery quality of different deepfakes varies: some have easily recognizable forgery clues, while others are highly realistic. Existing works often train detectors on a mix of deepfakes with varying forgery qualities, potentially leading detectors to short-cut the easy-to-spot artifacts from low-quality forgery samples, thereby hurting generalization performance. To tackle this issue, we propose a novel quality-centric framework for generic deepfake detection, which is composed of a Quality Evaluator, a low-quality data enhancement module, and a learning pacing strategy that explicitly incorporates forgery quality into the training process. Our framework is inspired by curriculum learning, which is designed to gradually enable the detector to learn more challenging deepfake samples, starting with easier samples and progressing to more realistic ones. We employ both static and dynamic assessments to assess the forgery quality, combining their scores to produce a final rating for each training sample. The rating score guides the selection of deepfake samples for training, with higher-rated samples having a higher probability of being chosen. Furthermore, we propose a novel frequency data augmentation method specifically designed for low-quality forgery samples, which helps to reduce obvious forgery traces and improve their overall realism. Extensive experiments demonstrate that our proposed framework can be applied plug-and-play to existing detection models and significantly enhance their generalization performance in detection.
1 Introduction
@@
Image synthesis techniques have experienced rapid advancements in recent years, providing a new way of entertainment [37, 16, 56]. Unfortunately, this technology is also broadly misused to spread false misinformation and even bring political influence. To prevent the abuse of deepfake, it is increasingly crucial to develop a reliable deepfake detector. Most earlier deepfake detectors [1, 39, 5, 18, 21, 23] achieve satisfactory and promising detection results in the within-domain evaluation scenario, where the distributions of the training and testing data are similar. However, when encountering cross-domain evaluation, with previously unseen forgery methods or data sources, the performance of these models will drop significantly [28, 36, 58]. Thus, the generalization ability of these prior works is still limited and underexplored.
@@
One reason for the generalization issue is the model’s shortcuts [15] on the easy-to-spot artifacts from the training samples with low forgery quality [53, 40, 46]. For instance, the “irregular swap" fake samples are frequently observed in FF++ (a common training set) but rarely in CDF-v2 (a common testing set). “Irregular swaps" was first noticed by [46], where face-swapping involves pairs with distinct facial attributes, such as differences in race or gender. These irregular swaps often severely distort the human facial structure, leading to lower-quality fake samples. As illustrated in Figure 1(a), we see that when “irregular swaps" occur (i.e., swapping a white face with a black face or a female face with a male face), the easy-to-recognize deepfake samples are generated. To verify the impact of these lower forgery quality fakes on the model’s generalization, we selectively choose samples with different forgery qualities for training the same detector (i.e., Xception [8]).
(a) Qualitatively assess the forgery quality from two views: swapping pairs (static) and model feedback (dynamic).
(b) Quantitative evaluation of model generalization performance across varying forgery quality levels.
Figure 1: Impact of forgery quality on deepfake detection. (a) Qualitative assessment of forgery quality from two perspectives: swapping pairs (static) and model feedback (dynamic); (b) Quantitative analysis of how different forgery quality levels affect model generalization performance.
Results in Figure 1(b) verify that these low static quality samples indeed hurt the model’s generalization, especially in the Wild [66] and DFDC-P [12] datasets, where simply training with higher-quality samples can obtain about 10% improvement in the generalization performance.
However, most existing methods often train detectors on a mix of qualities of deepfake data, which might inevitably struggle with this issue (similar to the “baseline" in Figure 1(b)), as they treat all data as the same and ignore the impact of different forgery qualities.
@@
These observations motivate us to treat training data with forgery qualities differently. To this end, we need to address three critical considerations first: Question 1: How can we assess the quality level of a sample? Question 2: How to deal with low-quality fake data? Question 3: How can we maximize the utilization of all types of data?
@@
For Question 1, we assess forgery quality from two perspectives—static (data-driven) and dynamic (model-driven)—as shown in Figure 1(a). From the static view, we focus on swapping pairs in face-swapping. Deepfakes generated from similar identity pairs tend to be more realistic. High-quality samples show minimal visual artifacts and natural facial structure, while low-quality samples exhibit obvious distortions. To quantify this, we use a face recognition model (ArcFace [10]) to compute the cosine similarity between fake and corresponding real images (with aligned backgrounds); a lower similarity indicates lower static quality. From the dynamic view, we evaluate how confidently the detection model predicts each fake during training. Lower prediction probabilities suggest harder-to-detect fakes, thus indicating higher dynamic quality (see Figure  1(a) for illustration). By combining both static and dynamic scores, we define a Forgery Quality Score (FQS) to guide sample selection [^1] Unless otherwise noted, low-quality and high-quality samples refer to those with low and high FQS, respectively..
For Question 2, instead of completely discarding low-quality deepfakes with obvious artifacts, we enhance their realism through a method called Frequency Data Augmentation (FreDA). FreDA improves low-quality samples by combining the low-frequency components from real faces (which preserve overall facial structure and semantic integrity) with the high-frequency components from fake faces (which contain more consistent and generalizable forgery cues). This fusion results in augmented samples with more natural facial structures and retained forgery signals, effectively transforming low-quality samples into more realistic, high-quality ones. By doing so, FreDA reduces non-generalizable artifacts and strengthens the training data for better generalization.
For Question 3, to make full use of training samples with varying quality, we adopt a curriculum learning strategy that guides the model from easy to hard examples. We first compute the Forgery Quality Score (FQS) for each sample and use it to control sampling during training—samples with higher FQS are more likely to enter the hard sample pool. As training progresses, we gradually reduce the number of training samples, shifting focus toward harder examples. Meanwhile, we apply FreDA to enhance low-quality samples, increasing their difficulty and diversifying the training set. This progressive training approach enables the model to benefit from the full spectrum of data quality.
Extensive experiments show it consistently improves both in-dataset and cross-dataset generalization across various deepfake detectors in a plug-and-play fashion. Notably, our method outperforms the baseline [8, 44, 34] by about 10% on average across several widely-used evaluation datasets, such as Celeb-DF [31] and DFDC [11].
Overall, our main contributions are as follows:
@@
  • We propose a quality assessment method combining both static and dynamic views to obtain the Forgery Quality Score (FQS), which allows us to rank the hardness of the training samples, thereby implementing the training sample selection.
  • we propose Frequency Data Augmentation (FreDA) to reduce the obvious forgery traces and thereby enhance the realism of the augmented low-quality data, rather than directly discarding them. FreDA can reduce the traces of forgery in forged samples. Using samples processed by FreDA enhances the model’s generalization.
@@
2 Related Works
@@
Most detection works at an early stage, mainly focus on hand-crafted features, e.g., eye-blinking [29], inconsistencies [60] of head poses [30], and other visual biological artifacts. With the rapid development of deep learning, data-driven-based detectors [39] have shown better performance than the conventional hand-crafted approaches. However, these approaches often suffer from poor generalization when there is a distribution shift between training and testing forgeries [36, 58, 55].
@@
Recent works have sought to improve generalization from different directions:
@@
1) Forgery Artifact Learning: Several solutions, such as disentanglement learning [57, 32], reconstruction learning [4, 47] and inconsistent learning [65, 49, 45, 61, 52, 13], have been proposed to learn general forgery artifacts. These methods aim to isolate the common characteristics of manipulated content, irrespective of the specific forgery technique employed. To illustrate, UCF [57] proposed a disentanglement-based framework to eliminate the overfitting of both forgery-irrelevant content and forgery-specific artifacts. CADDM [13] used the identity inconsistency between the inner and outer faces for detection. Additionally, However, these approaches mostly operate at the spatial level but overlook frequency-level artifacts, which have been demonstrated to be crucial for detection [38]. To this end, several studies have incorporated frequency information to enhance the performance of detectors [14, 38, 33, 50, 41]. Additionally, several recent works [24, 62, 48, 54] have proposed to leverage both local and global interactions to more accurately capture forgery traces. Furthermore, [43, 35, 25] leveraged contrastive learning to improve the detection performance of deepfake detectors.
@@
2) Forgery Augmentation: Forgery augmentation techniques have proven to be highly effective in enhancing the generalization ability of deepfake detectors by encouraging the model to learn generic forgery artifacts. Pioneering works such as BI (Face X-Ray) [28] synthesized blended faces by replicating blending artifacts between pairs of pristine images with similar facial landmarks. I2G [63] employed a similar strategy, leveraging pair-wise self-consistency learning to identify inconsistencies within generated fake images. Later efforts, like SBI [40], focused on self-blending, selecting the same source and target faces to train models, while ProDet [7] explored the way to fuse augmented forgery data with actual deepfake data more effectively. More recently, [51] and [59] have explored video-level blending for data augmentation, achieving generalizable results in detecting video deepfakes.
@@
3) Quality-aware deepfake detection: Another notable direction is to train the detector with different image qualities (combining both low-quality compression images and high-quality non-compression ones), aiming to learn the quality-irrelevant forgery artifacts [26]. Recent works [22, 42] proposed to investigate the correlation between the detection performance and the image quality. Additionally, [27] proposed a super-resolution model for improving low-quality deepfake detection. However, most existing works define the “quality" as the image quality, e.g., whether the image is blurry or not. Although our framework is also quality-based, our focus is more on the forgery quality, i.e., whether the deepfake image is realistic or not, rather than the image quality. In our framework, we propose to assess the forgery quality from two distinct aspects: swapping pairs and model feedback. We then design a curriculum-learning-based method to leverage samples with varying forgery qualities dynamically for training, guiding the detector to learn the samples from easy to hard gradually.
3 Method
In this section, we introduce the proposed quality-centric framework. Our framework is inspired by curriculum learning that explicitly encourages the detection model to learn forgery artifacts gradually from easy to hard based on FQS. It emphasizes three key components: firstly, the evaluation of data quality; secondly, the data augmentation on low-quality data; and thirdly, the progressive use of quality-based data during training. To achieve this, we carefully design three specified modules to implement the framework: 1) FQS is proposed as a metric to assess sample hardness during model training. It consists of two main components: static scores and dynamic scores. Static score (related to data) provides initial guidance at the start of model training. The weight of the static score gradually decreases as the training progresses. Dynamic score (related to the model) is continuously updated based on the model’s feedback, to better screen samples throughout the training process; 2) A FreDA module is proposed to enhance the forgery quality (“realism") of the augmented low-quality fake samples; 3) A pacing function is presented to control the pace of presenting data from easy to hard according to FQS. Each module will be detailed subsequently.